# A Survey of Conversational Search

### v1

@@ 6.2 User-centric Conversational Search
Conversational search systems are effectively employed across various scenarios to support users in seeking information. Instead of mainly focusing on specific predefined model abilities, such as retrieving world knowledge, reasoning across various candidate evidence, etc, some existing studies [150] concentrate on the perspective of user satisfaction, i.e., to improve users’ satisfaction by addressing their information needs under distinct intents. This is non-trivial because explicitly requesting satisfaction feedback from users is challenging and the approaches employed in traditional IR have not been thoroughly investigated in the context of conversational search. To facilitate the studies on user-centric conversational search, some benchmarks [150, 263] have been proposed recently for evaluating the system from a user-centric perspective and investigate how to build up proactive conversational search systems [151] together with human-centered agents [152]. Besides, some existing studies [153, 80] focus on mixed-initiative interactions, aiming to provide a better experience with multiple feedback to the user. Thus, maintaining a user-centric system is essential to ensure these techniques continue to serve users’ interests in future development.
Another research line attempts to facilitate user-centric conversational systems via personalized search, which requires satisfying users’ information needs and considering their backgrounds when searching candidate documents [10]. Specifically, the search results from the systems should yield different results for the same query according to the user profiles. The primary challenge is the issue of data scarcity in model development. To this end, Joko et al. [154] construct a large-scale personalized multi-session conversational search dataset synthetically by LLMs. Mo et al. [155] explore how to integrate query rewriting with personalized elements in a zero-shot manner by LLMs. Some recent studies aim to define concrete scenarios and evaluate the personalized ability of the systems [156, 10]. With the development of the data resource, more sophisticated personalized systems are expected.
6.3 Data Resources
The main bottleneck for developing domain-specific conversational search systems lies in data scarcity, which requires the knowledge resources to be injected into the models or serve as the external base. Besides, the available benchmark is crucial to evaluate the system’s specific ability corresponding to the requirements of various domains. To the best of our knowledge, we present the statistical detail of the data resources from each domain in Table I, which reflect the development of each special domain to some extent. The data resources provide the feasibility to train and evaluate the more sophisticated models in the future.
@@ 6.3 Data Resources
TREC iKAT [10] 2023 User-Centric Personalized Search en Text, User -
@@
TABLE I: An overview and statistic detail of the domain-specific conversational search data resource. $\ddagger$ denotes several subsets associated with different sub-tasks are contained.
@@ 6.4 Limitations and Discussion
Data Curation. The challenge in domain-specific tasks mainly lies in the lack of a comprehensive document repository, which constrains the feasibility for model training and evaluation. Compared with the general domain, domain-specific data are much higher cost to obtain due to the requirement of domain knowledge for the annotator. Thus, large-scale resources are desirable for domain-specific development and the automatically data generation techniques, e.g., LLM-aided, could be a potential avenue.
Domain-Specific IR. The complex terminology used in specific domains increases the difficulty of ensuring the matching between the retriever and query semantics, especially involved with local cultural, historical, political, and regional backgrounds. This could become extremely severe when the user lacks of expert knowledge and expresses the terminology in different ways to query. One potential solution is to leverage the domain-specific LLMs to help reformulate queries and the further development of domain-originated search engines by considering multi-factors are important.
@@ 7 Benchmark and Evaluation
A complete and intelligent conversational search system should be integrated with multiple components, including query reformulation, search clarification, conversational retrieval and response generation, etc. With the flexible interactive interface of the conversational search system, the information return form could depend on specific queries, which could be a rank list of the relevant documents, a summarization of them, a retrieval-augmented generated response, etc. The evaluation of existing conversational search systems can be divided into two categories [150]: 1) based on ranking metrics (e.g., MRR, NDCG, Recall), or 2) based on the semantic similarity metrics (e.g., F1, BLUE, and ROUGE), where the former is usually applied to retrieval tasks while the latter is commonly used with generation tasks. In this section, we first present different aspects of evaluation in existing studies with summarized available benchmarks. Then, we discuss the overlooked issues on retrieval-based and generation-based evaluation and propose some potential directions for improvement.
@@ 7.1 Evaluation Overview
The overview of the existing conversational search benchmarks is presented in Table II, where each benchmark corresponds to a specific setting, e.g., whether requiring to generate a final answer (RAG), involving a topic-switch (TS) phenomenon, or multiple reaction goal (Mixed-initiative) in a conversation. We can find that in the era of the large language model (after 2021), the user-iterative scenario and the RAG task have become more important and popular. Thus, different from the single-turn setting that most focuses on the accuracy of the results, it is crucial to evaluate the retrieval and generation task with various aspects from a conversational perspective.
@@ 7.1.1 Retrieval-based Evaluation
Following the traditional single-turn ad-hoc retrieval, existing studies on conversational retrieval commonly treat every query turn within a conversational session as a single query with corresponding relevance judgment as ground truth. This is because the available conversational search datasets are constructed inheriting from the single-turn datasets. In addition to the traditional ad-hoc search evaluation and focus on conversational features, Fu et al. [157] examine whether the same pattern of the assessment in ad-hoc search exists in conversational search based on both the user satisfaction and the observed performance of the search. Li et al. [158] and Siblini et al. [159] argue that the existing benchmarks using ground-truth answers provided in historical context are unreliable to evaluate the developed models regarding real human-machine conversations. To simulate an open-ended environment with unpredictable conversation trajectories for the practical situation, a series of studies leverage user simulators to provide feedback to the system’s responses. Early works [160, 161] study how to estimate the effectiveness of conversational search based on user simulation. Then, Wang et al. [162] simulate the risk of clarifying user questions and develop a risk-aware control model in terms of the search experience. To further facilitate the simulation of user feedback for conversational search, some recent studies [163, 164, 165, 166] develop user simulator-based frameworks to build up more user-system interaction as mixed-initiative to improve the conversational search systems via specific user feedback. Besides, it can also be used to evaluate mixed-initiative conversational search systems [167]. The goal of mixed-initiative systems is to enrich the interaction pattern, which can take initiative at any moment in the conversation and offer the user various questions or suggestions to address the information needs eventually. To evaluate the mixed-initiative aspect, Aliannejadi et al. [206] analyze the potential for more engaging conversational searches and the difficulty of assessing them due to the broader interaction scope. To this end, Sekulic et al. [167] achieve automatic evaluation by proposing a conversational user simulator. Besides, some recent studies [168, 169] focus on analyzing the prediction of system initiative and user action, which provide a new perspective for evaluating mixed-initiative systems.
@@ 7.1.2 Generation-based Evaluation
The flexible user-system interaction provides the chance to address complex information needs by multi-turn querying, which means the requirement of the user might not only be a rank list but also a concise answer from the retrieved results. Such a paradigm has become significantly more popular in the era of LLMs, making generation-based evaluation essential. However, conventional metrics such as F1 score, BLEU score, and exact match (EM) might not fully capture the quality of the generation. On one hand, some reasonable and correctly generated answers might not be annotated as the evaluation gold standard and fail in the traditional token-level metrics. On the other hand, the LLMs can yield accurate responses despite imperfect retrievers or hallucinations with semantic relevant retrieved text [253], raising the issue of evaluating the system by taking into account both retrieval and generation. Aim to combine the top retrieved documents into a comprehensive, pertinent, and concise response, Lajewska et al. [171] collect snippet-level answer annotations for existing conversational search benchmarks to fill the evaluation gap between retrieval and generation. To improve the correlation of retrieval effectiveness with the performance of the generation procedure, Salemi et al. [170] propose the eRAG metric, which evaluates the utilization of each retrieved document based on the ground truth labels of downstream tasks. Nevertheless, the evaluation gap remains between the aspect of IR and RAG, i.e., an IR-relevant document might not be effective for the downstream RAG task [252]. Thus, how to assess the performance of retrievers within LLM-based generators and better align the retriever-generator with the overall conversational search system is still an open challenge, which requires further exploration [172].
@@ 7.2 Limitations and Discussion
Although existing conversational search models have achieved significant development, the convinced evaluation of these systems is still an open question. The main challenge lies in how to evaluate them with several aspects which will be discussed as follows.
@@ 8 Conclusion and Future Directions
We expect our effort in this survey can illustrate the required landscape of modern conversational search and provide a strong foundation for future research in this rapidly evolving field. Aiming for more intelligent conversational search systems for future development, the overall suggestions are summarized as follows:
@@
  • Intelligent Decision Making. The appropriate system returns as the interactions with the user are the guarantee for a good search experience along with the conversation diving in. An intelligent decision maker is desirable to determine whether requiring clarification, external retrieval, additional response generation, etc. An alternative is to leverage LLM agents to complete complex decision tasks with tool-call, advanced memory, reasoning, and planning capabilities. The challenges lie in ensuring relevance and accuracy, mitigating biases, providing real-time responses, etc.
@@
  • Faithful and Abundant Resources. Faithfulness is crucial for search systems, which is an open challenge for generative-based systems. The key issue is to ensure the generated content are reliable response for the user, by indicating resources for knowledge attribution. Besides, to make resources abundant, e.g., integrating multi-modality information can also increase credibility and improve the results diversification of the interactive systems.
@@
  • Proactive and Personalized Conversation. The existing conversational search systems mostly perform in reactive ways, which respond to the user for every query turn until the user terminates the search session. A more optimal interaction design is acting proactively, where the system can decide and engage in the conversation. This is also the vision of mixed-initiative interactions, while there is a lack of data to train and evaluate the systems. Another important element is to offer personalized search, which requires the conversation systems to consider appropriate aspects from the user profile for each query to help generate personalized content as the final response.
@@
  • Toward Practical Evaluation. The existing datasets for evaluation are based on a strong assumption that all the previous turns are answered correctly when dealing with the current turn. This is far from the practical situation because the changed user behaviors will influence the later conversation flows. To address this open challenge, some efforts aim to develop user simulators to mimic real-world scenarios for practical evaluation. Further exploration for benchmark developments and new evaluation paradigms is desirable for the research community.
@@
  • Informative Returns. With the development of conversational search systems, the returning information might not limited to clarification questions or generated responses. It might emphasize more on informative forms and proactive interaction. For example, the conversational system could navigate the user toward different directions, e.g., calling external tools to tackle the request [272], planning a multi-step strategy as thought chains to solve the information needs [273], and collaborating with the environment to conduct agentic information retrieval [274]. The key idea is to enlarge the available resources and capacity to help users reflect during the conversation and eventually achieve their information needs.

### v2

@@ 6.2 User-centric Conversational Search
Conversational search systems are effectively employed across various scenarios to support users in seeking information. Instead of mainly focusing on specific predefined model abilities, such as retrieving world knowledge, reasoning across various candidate evidence, etc, some existing studies [175] concentrate on the perspective of user satisfaction, i.e., to improve users’ satisfaction by addressing their information needs under distinct intents. This is non-trivial because explicitly requesting satisfaction feedback from users is challenging and the approaches employed in traditional IR have not been thoroughly investigated in the context of conversational search. To facilitate the studies on user-centric conversational search, some benchmarks [175, 200] have been proposed recently for evaluating the system from a user-centric perspective and investigate how to build up proactive conversational search systems [176, 182] together with human-centered agents [177]. Besides, some existing studies [178, 78] focus on mixed-initiative interactions, aiming to provide a better experience with multiple feedback to the user. Thus, maintaining a user-centric system is essential to ensure these techniques continue to serve users’ interests in future development.
@@
Qilin [287] 2025 Multi-modal Search zh Text, Image, Product - 5.7k
@@
TREC iKAT 2023 [12] 2023 User-Centric Personalized Search en Text, User - 176
TREC iKAT 2024 [231] 2024 User-Centric Personalized Search en Text, User - 103
@@
TABLE I: An overview and statistical details of the domain-specific conversational search data resource. $\ddagger$ denotes several subsets associated with different sub-tasks are contained.
Another research line attempts to facilitate user-centric conversational systems via personalized search, which requires satisfying users’ information needs and considering their backgrounds when searching candidate documents [12, 231]. Specifically, the search results from the systems should yield different results for the same query according to the user profiles. The primary challenge is the issue of data scarcity in model development. To this end, Joko et al. [179] construct a large-scale personalized multi-session conversational search dataset synthetically by LLMs. Mo et al. [180] explore how to integrate query rewriting with personalized elements in a zero-shot manner by LLMs. Some recent studies aim to define concrete scenarios and evaluate the personalized ability of the systems [181, 12]. With the development of the data resource, more sophisticated personalized systems are expected.
6.2.1 Ethical Considerations and Bias
A user-centric conversational search system should take the aspect of ethical considerations and potential bias into account. Since the LLMs are trained on vast web data, LLM-based conversational systems could inherit and propagate societal biases related to gender, race, socioeconomic status, and other attributes [288]. As a result, the skewed or unfair information presentation might be presented in search results or generated responses. This bias would be amplified, especially when it is connected with user profiles or interaction logs [289]. Besides, conversational interactions can involve users revealing sensitive personal information. As conversational search systems will be widely deployed in the future and used by daily users more frequently, systems must employ robust privacy-preserving techniques [290] for data handling, storage, and model training, especially when personalization is involved.
6.3 Data Resources
The main bottleneck for developing domain-specific conversational search systems lies in data scarcity, which requires the knowledge resources to be injected into the models or serve as the external base. Besides, the available benchmark is crucial to evaluate the system’s specific ability corresponding to the requirements of various domains. To the best of our knowledge, we present the statistical detail of the data resources from each domain in Table I, which reflect the development of each special domain and provide the feasibility to train and evaluate the more sophisticated models.
@@ 6.4 Limitations and Discussion
Data Curation. The challenge in domain-specific tasks mainly lies in the lack of a comprehensive document repository, which constrains the feasibility for model training and evaluation. Compared with the general domain, domain-specific data are much higher cost to obtain due to the requirement of domain knowledge for the annotator. Thus, large-scale resources are desirable for domain-specific development and the automatically data generation techniques, e.g., LLM-aided, could be a potential avenue [291, 292].
Domain-Specific IR. The complex terminology used in specific domains increases the difficulty of matching the retriever and query semantics, especially when involved with local cultural, historical, political, and regional backgrounds. This could become extremely severe when the user lacks of expert knowledge and expresses the terminology in different ways to query. One potential solution is to leverage the domain-specific LLMs to help reformulate queries. Besides, the further development of domain-originated search engines by considering multi-factors are important.
@@ 7 Benchmark and Evaluation
A complete and intelligent conversational search system should be integrated with multiple components, including query reformulation, search clarification, conversational retrieval and response generation, etc. With the flexible interactive interface of the conversational search system, the information return form could depend on specific queries, which could be a rank list of the relevant documents, a summarization of them, a retrieval-augmented generated response, etc. The evaluation of existing conversational search systems can be divided into two categories [175]: 1) based on ranking metrics (e.g., MRR, NDCG, Recall), or 2) based on the semantic similarity metrics (e.g., F1, BLUE, and ROUGE), where the former is usually applied to retrieval tasks while the latter is commonly used with generation tasks. In this section, we first present different aspects of evaluation in existing studies with summarized available benchmarks. Then, we discuss the overlooked issues in retrieval-based and generation-based evaluation and propose some potential directions for improvement.
@@ 7.1 Evaluation Overview
The overview of the existing conversational search benchmarks is presented in Table II, where each benchmark corresponds to a specific setting, e.g., whether requiring to generate a final answer (RAG), involving a topic-switch (TS) phenomenon, or multiple reaction goal (Mixed-initiative) in a conversation. We can find that in the era of the large language model (after 2021), the user-iterative scenario and the RAG task have become more important and popular. Thus, different from the single-turn setting that most focuses on the accuracy of the results, it is crucial to evaluate the retrieval and generation task with various aspects from a conversational perspective. Besides, the widespread exposure to powerful LLMs has significantly shifted user expectations and behaviors, e.g., from one-shot search to multi-turn interactions, which raises the new requirement for developing evaluation protocols from the aspect of users.
@@ 7.1.1 Retrieval-based Evaluation
Following the traditional single-turn ad-hoc retrieval, existing studies on conversational retrieval commonly treat every query turn within a conversational session as a single query with the corresponding relevance judgment as ground truth. This is because the available conversational search datasets are constructed by inheriting from the single-turn datasets. In addition to the traditional ad-hoc search evaluation and focus on conversational features, Fu et al. [184] examine whether the same pattern of the assessment in ad-hoc search exists in conversational search based on both the user satisfaction and the observed performance of the search. Li et al. [185] and Siblini et al. [186] argue that the existing benchmarks using ground-truth answers provided in historical context are unreliable to evaluate the developed models regarding real human-machine conversations. To simulate an open-ended environment with unpredictable conversation trajectories for the practical situation, a series of studies leverage user simulators to provide feedback to the system’s responses. Early works [187, 188] study how to estimate the effectiveness of conversational search based on user simulation. Then, Wang et al. [189] simulate the risk of clarifying user questions and develop a risk-aware control model in terms of the search experience. To further facilitate the simulation of user feedback for conversational search, some recent studies [190, 191, 192, 193] develop user simulator-based frameworks to build up more user-system interaction as mixed-initiative to improve the conversational search systems via specific user feedback. Besides, it can also be used to evaluate mixed-initiative conversational search systems [194]. The goal of mixed-initiative systems is to enrich the interaction pattern, which can take initiative at any moment in the conversation and offer the user various questions or suggestions to address the information needs eventually. To evaluate the mixed-initiative aspect, Aliannejadi et al. [240] analyze the potential for more engaging conversational searches and the difficulty of assessing them due to the broader interaction scope. To this end, Sekulic et al. [194] achieve automatic evaluation by proposing a conversational user simulator. Besides, some recent studies [195, 196] focus on analyzing the prediction of system initiative and user action, which provide a new perspective for evaluating mixed-initiative systems.
@@ 7.1.2 Generation-based Evaluation
The flexible user-system interaction provides the chance to address complex information needs by multi-turn querying, which means the requirement of the user might not only be a rank list but also a concise answer from the retrieved results. Such a paradigm has become significantly more popular in the era of LLMs, making generation-based evaluation essential. However, conventional metrics such as F1, BLEU score, and exact match (EM) might not fully capture the quality of the generation. On one hand, some reasonable and correctly generated answers might not be annotated as the evaluation gold standard and fail in the traditional token-level metrics. On the other hand, the LLMs can yield accurate responses despite imperfect retrievers or hallucinations with semantic relevant retrieved text [267], raising the issue of evaluating the system by taking into account both retrieval and generation. Aim to combine the top retrieved documents into a comprehensive, pertinent, and concise response, Lajewska et al. [198] collect snippet-level answer annotations for existing conversational search benchmarks to fill the evaluation gap between retrieval and generation. To improve the correlation of retrieval effectiveness with the performance of the generation procedure, Salemi et al. [197] propose the eRAG metric, which evaluates the utilization of each retrieved document based on the ground truth labels of downstream tasks. Nevertheless, the evaluation gap remains between the aspect of IR and RAG, i.e., an IR-relevant document might not be effective for the downstream RAG task [266]. Thus, how to assess the performance of retrievers within LLM-based generators and better align the retriever-generator with the overall conversational search system is still an open challenge, which requires further exploration [199].
7.1.3 User-based Evaluation
Except for delving into specific metrics for accuracy, it is necessary to consider the practical impact, such as user satisfaction between traditional search interfaces and conversational search interfaces. The property of mixed-initiative of conversational search systems and their interfaces provides potential for higher user engagement and satisfaction, particularly for complex, exploratory information needs, which is inapplicable in traditional one-shot search. The multi-turn interactions can help refine queries and synthesize information effectively, which provides a different search experience. Thus, evaluating conversational systems with careful consideration of user experience and satisfaction within realistic interaction contexts is important. The previous studies in human-computer interaction [293, 280] and interactive information retrieval [294] aim to investigate how to maximize user utility [295] by emphasizing more on qualitative user studies, understanding cognitive load, interaction friction, and the nuances of dialogue success beyond simple definitions of relevance in search. In the era of LLMs now, users often expect more direct and synthesized answers, engage in longer and more exploratory dialogues, and may deploy the conversational intelligent systems for tasks beyond simple fact retrieval, such as task completion or content creation. Thus, the users might have higher expectations regarding the naturalness and coherence of the conversation, which consequently results in the requirement to evaluate user satisfaction or conduct user studies. Some recent studies aim to develop user-based evaluation standards from the aspects of user satisfaction in the task completion [200, 296], the interactions in the command-and-control paradigm [201], and detect-then-mitigate cognitive bias [202].
@@ 7.2 Limitations and Discussion
Although existing conversational search models have achieved significant development, the convincing evaluation of these systems is still an open question. The main challenge lies in how to evaluate them with several aspects as follows.
@@
User Satisfaction Evaluation in Conversation and Interaction. The general evaluation metrics focused solely on document relevance or turn-level answer accuracy are insufficient for modern conversational search systems. The evaluation toward holistic, task-based, user-centric metrics in the context of LLM-based systems could be connected with the findings from decades of human-computer interaction and interactive information retrieval research [298], especially user behaviors and expectations could be involved [299]. Besides, compared to traditional ad-hoc search, users might notice conversational systems are slower for simple fact retrieval, producing verbose responses, or have difficulties adapting to a multiple-turn-based interaction model compared to the familiar keyword-and-scan approach of traditional search. The higher latency, trustworthiness in LLM-generated answers, and the system’s ability to gracefully handle errors significantly impact user satisfaction and adoption rates. Thus, developing more meaningful evaluation protocols for current powerful, but complex, LLM-based conversational search systems is promising and desirable.
@@ 8 Conclusion and Future Directions
We expect our effort in this survey can illustrate the required landscape of modern conversational search and provide a strong foundation for future research in this rapidly evolving field. Aiming for more intelligent conversational search systems for future development, the overall suggestions and research questions are summarized as follows:
@@
Intelligent Decision Making. How to enable conversational search systems to robustly determine the optimal action, e.g., clarify, retrieve, generate, invoke an external tool, at each conversational turn, adapting dynamically based on dialogue state, user intent, and system confidence? The appropriate system returns as the interactions with the user are the guarantee for a good search experience, along with the conversation diving in. An alternative is to leverage LLM agents to complete complex decision tasks with tool-call, advanced memory, reasoning, and planning capabilities. The challenges lie in ensuring relevance and accuracy, mitigating biases, providing real-time responses, etc.
@@
Faithful and Abundant Resources. How to develop reliable knowledge attribution methods for multi-source conversational generation to mitigate hallucination? Faithfulness is crucial for search systems, which is an open challenge for generative-based systems. The key issue is to ensure the generated content are reliable response for the user, by indicating resources for knowledge attribution. Besides, to make resources abundant, e.g., integrating multi-modality information, can also increase credibility and improve the results diversification of the interactive systems.
@@
Proactive and Personalized Conversation. How can we enable the conversational search systems to proactively interact with the users to improve the search experience? The existing conversational search systems mostly perform in reactive ways, which respond to the user for every query turn until the user terminates the search session. A more optimal interaction design is acting proactively, where the system can decide and engage in the conversation. This is also the vision of mixed-initiative interactions, while there is a lack of data to train and evaluate the systems. Another important element is to offer personalized search, which requires the conversation systems to consider appropriate aspects from the user profile for each query to help generate personalized content as the final response.
@@
Toward Practical Evaluation. How to develop high-fidelity user simulations that simulate user behavior, including error recovery, adaptation, and evolving information goals, for practical system evaluation? The existing datasets for evaluation are based on a strong assumption that all the previous turns are answered correctly when dealing with the current turn. This is far from the practical situation because the changed user behaviors will influence the later conversation flows. To address this open challenge, some efforts aim to develop user simulators to mimic real-world scenarios for practical evaluation. Besides, how to develop methodologies and comprehensive benchmarks to capture and evaluate long-term conversational success, task completion, information gain, and user satisfaction for practical conversational search scenarios is also valuable and promising. Further exploration for benchmark developments and new evaluation paradigms is desirable for the research community.
@@
Informative Returns. How to enrich the information within the returning response? With the development of conversational search systems, the returning information might not be limited to clarification questions or generated responses. It might emphasize more on informative forms and proactive interaction. The system could navigate the user toward different directions, e.g., calling external tools to tackle the request [300], planning a multi-step strategy as thought chains to decompose the request [301], and collaborating with the environment to conduct agentic information retrieval [302]. The key idea is to enlarge the available resources and capacity to help users reflect during the conversation and eventually achieve their information needs.
New Paradigms of Conversational Search. How to integrate the developed capabilities of LLMs with the search engine toward a new paradigm of conversational search? Given recent advances with LLMs and their adoption into day-to-day activities, the research community needs to reconsider what we are ultimately trying to achieve with conversational search. Understanding these goals would help contextualize how prior works attempted to solve the technical problems within their limitations, informing the algorithms, metrics, and exploration designs used. A new paradigm of conversational search by both inherits from the previous studies and considers the impacts of recent advances is desirable.